A man walks at a pace of 3 miles per hour. After 2 hours he stops for a rest. Let t be time AFTER the rest when he continues on his walk. Let d be the TOTAL distance traveled since he started. Write an equation linking d and t.
Answer by ptaylor(1894)


Distance(d) equals Rate(r) times Time(t) or d=rt; r=d/t and t=d/r
After 2 hours the man has walked a distance of 3*2=6 mi
When the man continues his walk he will walk a distance of 3t mi
So:
Total distance:
d=6+3t---------------------------ans


Hope this helps---ptaylor